40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2148     Entry time: Tue Oct 27 01:45:02 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: MZ 

Tonight we also encountered a large peak in the frequency noise around 485 Hz. Changing the MZ lock point (the spot in the PZT range) solved this.

This again tonight.

It hindered the initial acquisition, and made the DD signal handoff fail repeatedly.
ELOG V3.1.3-